OpenStudy (anonymous):

Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 Equation 2 |5x − 6| = −41 |7x + 13| = 27 Equation 1 has more solutions than equation 2. Equation 1 and Equation 2 have the same number of solutions. Equation 2 has more solutions than Equation 1. The number of solutions cannot be determined.

OpenStudy (nurali):

|5x − 6| = −41 |7x + 13| = 27 Ans: Equation 1 5x - 6 = -41 5x = -41 + 6 5x = -35 5x/5 = -35/5 x = -7 Equation 2 -27 = |7x + 13| = 27 -27 = 7x + 13 -27 - 13 = 7x -40 = 7x -40/7 = x -5.7143 = x 7x + 13 = 27 7x = 27 - 13 = 14 7x/7 = 14/7 x = 2 x = -40/7, 2 Comparing, Equation 1 has only 1 solution While Equation 2 has 2 solutions. Thus, [ Equation 2 has more solutions than Equation 1 ]

OpenStudy (anonymous):

you are my hero lol